OpenStudy (anonymous):

A wheel rotates counterclockwise at a constant rate of once every 10 seconds. The center of the wheel is the origin and its radius is 1 foot. A point P is marked on the wheel and at time t=0sec it has coordinates (1,0). How far does P travel in 1 minute? If an angle formed by the positive x axis and a line from the origin through P, how many degrees does the point rotate in 1 minute? At 1.25 sec, what are the coordinates of P?

OpenStudy (mathmath333):



OpenStudy (mathmath333):

wheel rotates counter clock wise once in 10 sec each hence the rotation in 1 min will be \[\frac{ 60}{ 10 }=6\] times

OpenStudy (mathmath333):

the circumference of cir\[C =2 \pi r\]cle \[=2 \pi *1\] \[=2 \]

OpenStudy (mathmath333):

circumference C\[C=2 \pi r\] \[= 2 \pi *1=2 \pi\]

OpenStudy (mathmath333):

as it traved 6 times in one min hence the distance traveled in 1 min will be \[= 6*2 \pi =12 \pi\]

OpenStudy (mathmath333):